Suppose the market demand and supply curves are represented by the following equations:

Qd = 110 – 20P

Qs = -10 + 40P

  1. Determine the equilibrium price and quantity. (Show your calculations)

  1. Illustrate these curves on a graph, labelling the curves, intercepts and equilibrium price and quantity.
Answer #1

a) SET D = S

110-20P = -10+40P

110+10 = 40P+20P

120 = 60P

P = 120/60 = 2

Q = 110-20(2) = 70

b) Maximum reservation price = 110/20 = 5.5

Minimum reservation price = 10/40 = 0.25